OpenStudy (anonymous):

Simplify in Radical Expresion \[\sqrt{216}\]

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

Sqrt216=sqrt(6*6*6)=sqrt((6)^2*6)=6sqrt6

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

you are welcome:)

OpenStudy (anonymous):

\[\sqrt{512}\] ?

OpenStudy (anonymous):

rad form

OpenStudy (anonymous):

Here, we can rewrite it as...sqrt(512)=sqrt(2*2*2*2*2*2*2*2*2)=(sqrt(2^2)*sqrt(2^2)*sqrt(2^2)sqrt2=2*2*2sqrt2=8sqrt2

OpenStudy (anonymous):

Thx again. i dont really pay attention in class.

OpenStudy (anonymous):

no problem:)

OpenStudy (anonymous):

whenever u will get the same kind of problem just try to factor it with least value 2 and then go in increasing order..

OpenStudy (anonymous):

if 2 is not sufficient then try 3 then 4 ......and go on..it will help you